[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Nov 01, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Tue Nov 1 13:46:40 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The problem with the NHC web server over the weekend turned out to be  
caused by an expired security certificate and was not related to the  
computer room AC problem. That issue is being resolved by NCO.

The problem with the LDM on the logger server occurred again last night,  
resulting is products not posting on the NHC web page. Matt again repaired  
the LDM feed and set up some redundancy so the web products are also  
supplied by the NOAAPort feed. The problem with logger is still being  
investigated.

The LX5 AWIPS workstation at the surface desk froze up. Marshall contacted  
NCF and they were able to resolve the problem with a reboot.
HSU:
See the TAFB discussion about web server and LDM issues preventing products  
from updating on the NHC web page.
Other items:
The LDM on the logger server mentioned above is fed by the socket feed,  
which is also known as "line 162" by the TOC. That line is monitored by  
NCO, and when our LDM fails, NCO gets an error message and contacts NHC for  
updates (usually TAFB during off hours). If TAFB or HSU gets one of these  
calls about line 162 from the tech control group, let them know NHC is  
still working on the problem and make a note in the HSU or TAFB shift logs.

Work in the media room is resuming today.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161101/4ad3e205/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list